OpenStudy (anonymous):

The function below has at least one rational zero. Use this fact to find all zeros of the function. g(x)=3x^4-8x^3+4x^2-x+2 If there is more than one zero, separate them with commas. Write exact values, not decimal approximations.

OpenStudy (anonymous):

I get rational zeros \[\pm1,\pm \frac{ 1 }{ 3 }, \pm2,\pm \frac{ 2 }{ 3 }\]

terenzreignz (terenzreignz):

These are the POSSIBLE rational zeros. You need to plug them in and check them...

OpenStudy (anonymous):

So I get 1,2, 11i, -11i is that correct?

terenzreignz (terenzreignz):

Hang on :D So, you're right about 1 and 2... how did you get 11i and -11i ? Serious question by the way... I'm terrible at division of polynomials o.O

OpenStudy (anonymous):

using synthetic division to find the real zeros I got: 1/6 (-1-sqrt11) and 1/6(-1+sqrt11)

OpenStudy (anonymous):

which is a complex root

terenzreignz (terenzreignz):

My calculator says you're right :D

OpenStudy (anonymous):

So should I include the complex root in my answer? 1,2, -1/6(11i), and 1/6(11i)

OpenStudy (anonymous):

what calculator? Can I buy it? :D

terenzreignz (terenzreignz):

Yes. They said all the zeros, not just the rational zeros. (or real zeros, for that matter) It's a Casio calculator. I'm sure it's available somewhere in your vicinity. I asked it to (LOL) find the roots of the equation 3x² + x + 1 And it jives with your answers :D

OpenStudy (anonymous):

I'm doing the "Aleks" assessment. Just don't know how to insert the answer so it will read it the right way.

terenzreignz (terenzreignz):

whoop-de-doo Can't help you in that regard, mate :D Cheers ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@amistre I just want to post this the right way. 1,2, \[\frac{ 1 }{ 6 }(-1\pm i \sqrt{11})\] Is that the right way to write the complex root or should I separate it,i.e. \[\frac{ 1 }{ 6 }(-1-i \sqrt{11}), \frac{ 1 }{ 6 } (-1+i \sqrt{11})\]

OpenStudy (anonymous):

@amistre64 , @Mertsj

OpenStudy (mertsj):

Don't they give instructions about the correct way to report answers?

OpenStudy (amistre64):

Use this fact to find \(\underline{ALL}\) zeros of the function. If it had stipulated all REAL zeros then we would omit the complex ones

OpenStudy (anonymous):

Gotcha!

OpenStudy (amistre64):

pfft, formating eh ... thats up to the program

OpenStudy (anonymous):

I know right!

OpenStudy (amistre64):

separate them with commas, so second option

OpenStudy (anonymous):

but is it typical to put the plus/minus sign into the answer or do you normally separate it?

OpenStudy (anonymous):

Got it- thank you guys!

OpenStudy (amistre64):

mathically, its the same .... programically ... it matters

OpenStudy (anonymous):

I see said the mathmatically blind man.

OpenStudy (anonymous):

You guys are the best- thanks again.